# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v2

@@ Abstract.
Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life. While essential in various scenarios, confinement within the home environment can profoundly influence psychological well-being and digital device usage. In this study, we delve into these effects, utilising the COVID-19 lockdown as a special case study to draw insights extending to various homebound situations. We conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks and analysed their emotions, well-being, social roles, and mobile usage behaviours. We extracted user activity from app usage records in an unsupervised manner, and experimental results revealed that app usage behaviours are effective indicators of emotional well-being in confined environments. Our research has great potential for developing supportive strategies and remote programs, not only for people facing similar medical isolation situations, but also for individuals in long-term home confinement, such as those with chronic illnesses, recovering from surgery, or adapting to permanent remote work arrangements.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXjournal: TOCHIccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
@@ 1. Introduction
Home confinement, defined as the restriction of individuals’ movements and activities within their homes (63), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (47; 17). Previous research indicated that home confinement may adversely affect mental well-being, leading to stress, anxiety, and depression (7; 78). For instance, Bartoszek et al. (7) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue. Understanding the impact of home confinement on mental well-being is essential for developing effective interventions to support individuals during prolonged periods of confinement.
Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (78; 30). However, these approaches can be time-consuming and labour-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (32; 71)). Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being unobtrusively (73; 21). For instance, Rashid et al. (73) modelled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al. (21) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices.
In home confinement scenarios, traditional mobile sensing methods, heavily relying on sensors like accelerometers and GPS to track physical behaviours, encounter unique challenges. These methods, which infer mental well-being from patterns such as visited locations, steps taken, and stationary periods (73), become less effective when individuals are confined at home. During home confinement, individuals’ interaction with their smartphones changes significantly. Rather than carrying their devices with them, people tend to leave their phones in one place. This shift in usage patterns diminishes the effectiveness of conventional metrics such as step counts and travel patterns, which are less indicative of mental well-being in a stationary context. Consequently, there is a shift towards analyzing more nuanced data types, such as application usage patterns, to better understand well-being under home confinement. Therefore, in this research, we focus on these finer-grained user activities to uncover deeper insights into the subtle changes in mental well-being during home confinement.
Another challenge in studying the in-situ mental well-being of individuals confined to their homes is establishing a natural yet controlled environment. Home confinement situations vary significantly among individuals, influenced by different reasons and individual differences, making effective modelling challenging. Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement. The strict measures in place have created similar confinement conditions for all participants due to policy mandates. Additionally, the COVID-19 home confinement has led to an increasing reliance on digital technologies (68), with strategies such as stay-at-home and work-from-home being recommended worldwide (74). Recent research indicates that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (82). This reliance on technology provides an excellent way to understand people’s mental well-being during home confinement.
In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during home confinement. Specifically, we focus on emotions and social roles, as these aspects significantly impact overall well-being (23; 65). To assess emotional states, we utilize the widely accepted Valence and Arousal dimensions within the Circumplex Model of Affect (79) to assess emotional states. Social roles, as defined by Role Theory (9), involve norms, expectations, rules and behaviours that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance. This, in turn, can have a significant impact on an individual’s overall well-being and mental health. Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behaviours. As a result, our research questions are:
  • RQ1. What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?
  • RQ2. Is it possible to unsupervisedly extract user activity from app usage records? If so, how are these activities related to emotions, productivity and interruptions?
  • RQ3. Can we model people’s emotion from mobile usage during home confinement? If so, what are the most significant types of features for prediction?
@@
To answer the research questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia. To answer RQ1, we analyzed patterns of perceived valence, arousal and social roles during COVID-19 home confinement using Experience Sampling Method (ESM) (8). We also investigated the influence of the COVID-19 lockdown on emotions, social roles, and work-life balance by examining the end-of-day survey data. To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day. Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviours for RQ3. All scripts for this study are open source. 111link to be shared after the double-blind review. In summary, our contributions are as below:
@@
  • We collected a comprehensive dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia. The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc. Their smartphone and desktop usage were recorded through an app running in the background. The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
@@
  • We analyzed participants’ perceived valence, arousal, and social roles during home confinement, and explored how lockdown affected well-being and work-life balance. We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
  • We proposed an unsupervised activity recognition method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records. This approach autonomously identified 16 distinct user activities in an unsupervised manner, and we then explored the correlations between these user activities and human well-being.
  • We conducted extensive experiments to predict human well-being in confined environments from smartphone usage behaviours, demonstrating the superior performance of our model. Additionally, we identified key factors contributing to well-being and discussed the primary predictors of human emotion during home confinement.
@@ 2. Related work
2.1. Impact of the Pandemic Restrictions on Human Well-being
Pandemics and epidemics have played an essential role throughout human history (87). In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world. Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (87; 20; 12), including post-traumatic stress symptoms (76), depression (35), stress (26), emotional disturbance (97), irritability(48), low mood (48), confusion (16), anger (56) and anxiety-induced insomnia (26). Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (42).
According to the World Health Organization  (94), the COVID-19 pandemic has severely impacted human well-being. The potential for physically sick or discomfort from COVID-19 is considered less of a concern than the psychological and social impact (38). A growing body of literature has explored the impact of COVID-19 on emotional and mental health (95; 22; 49; 46). Daly et al.  (22) explored the effects of COVID-19 on mental health in the UK and found a significant increase in observed mental health problems. Fountoulakis et al. (28) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased by 10.40%. Khubchandani et al. (45) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after the initial COVID-19 lockdown in the USA. The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019. Killgore et al. (46) found that loneliness has increased during the COVID-19 pandemic, especially for those under lockdown restrictions.
Despite the existing research on the impact of pandemic restrictions on human well-being, much of it is limited to specific regions and conducted during different phases of varying degrees of restrictive measures. This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being. In light of this, RQ1 focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage IV) were implemented.
2.2. Impact of the Pandemic Restrictions on Smartphone Usage
The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (89; 27; 4). Twitter even made an announcement to make working remotely a permanent option (13). As people perform their duties by staying at home or working from home, they increasingly turn to digital devices to communicate, interact and conduct job responsibilities (74). Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (82; 68). A recent global study (82) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (74).
Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (74; 14; 5). Ratan et al. (74) pointed out that problematic smartphone usage affects the ‘reward system’ of the brain and subsequent harms will persist in the form of various emotional disorders even after the lockdown effects are lifted. Arora et al. (5) explored the problematic phone use and its impact on mental health during the COVID-19 lockdown, and found that ‘sadness’ and ‘anger’ were the dominant emotions in addictive Twitter users, while ‘valence’ was the dominating emotion among non-addictive users. Busch et al. (14) revealed that excessive mobile phone use may severely impact people’s mental health and human well-being. Especially during the COVID-19 lockdown, people’s mobile phone usage differs than normal, which increases the risk of overusing smartphones and affects their mental health (6).
2.3. Investigating User Activities from App Sequences
By analysing app sequences, researchers have gained insights into various aspects of individuals’ daily behaviours, including social roles  (2) and interruptiblity (66; 67). Deploying the approach ‘Application as a Sensor’, Anderson et al.  (2) found that app sequences on smartphones are correlated with individuals’ private and work social roles. Similarly, Okoshi et al.  (66; 67) utilized app sequence data and machine learning techniques to predict breakpoints in activity for intelligent notification delivery. These studies focus on identifying breakpoints between apps to understand interruptibility. However, no previous research has used breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period. However, those approaches fail to capture the nuances and transitions between different app activities, limiting understanding of user behaviours. Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data. Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information about user behaviours and mental well-being. With this in mind, RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
@@ 2.3. Investigating User Activities from App Sequences
2.4. Inferring Human Well-being Using Mobile Sensing Technologies
The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (55) and psychological stages, e.g., valence (41), mood (60; 98; 59), engagement (33; 39; 31), social anxiety (21; 73), depression (91; 88), etc. Most studies integrate data from various sources for building comprehensive prediction models. For instance, Gao et al. (33) predicted emotional, cognitive and behavioural student learning engagement in an Australian high school using wearable sensors and environmental sensors. Wang et al. (91) measured depression of undergraduate students at Dartmouth College using passive mobile sensing data and wearable data. Jaques et al. (41) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behaviour data collected from mobile devices. Morshed et al. (60) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behaviour, e.g., physical activity, communication patterns, and location. For example, Hintze et al. (37) analyzed mobile device usage characteristics such as session length, interaction frequency and daily usage in locked/unlocked state with respect to location context and diurnal pattern, to gain insights into how people use their smartphones and tablets. Madan (55) et al. designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (73; 88; 21), mood (98; 59), emotion (86), social roles (2) and interruptibility (66; 67). Zhang et al. (98) proposed an Android app to detect compound emotion utilizing smartphone usage data and Meegahapola et al. (59) explored the effect of geographical diversity on mood inference. Tlachac et al. (88) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine-learning models to predict anxiety and depression. Tag et al. (86) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings.
@@ 2.4. Inferring Human Well-being Using Mobile Sensing Technologies
While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement. Unlike daily life situations, individuals’ behaviours may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working. Although a few studies (88; 64) have examined psychological states using mobile sensing during the COVID-19 pandemic, these works largely focus on broader environments rather than strictly home confinement. Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, ageing or disability-related limited mobility. Therefore, we aim to address RQ3 by exploring the possibility of modelling emotions during home confinement and identifying significant features for prediction.
3. Data Collection
We conducted a field study in Melbourne, Australia, for three weeks, from 24 August 2020 to 13 September 2020. The study has been approved by the Human Research Ethics Committee at the authors’ university. Prior to the data collection, we obtained informed consent from all participants, ensuring that they were aware of the study’s purpose and that the dataset would be publicly released. The details about the background, participants, and collected data will be described in this section.
3.1. Background

### v3

@@ Abstract.
Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life. While essential in various scenarios, confinement within the home environment can profoundly influence mental well-being and digital device usage. Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of prolonged home confinement on mobile device usage. We conducted an in-situ study with 32 participants, analyzing three weeks of mobile usage data to assess emotional well-being and social dynamics in restricted environments. Our findings reveal that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation. This study highlights the potential for developing targeted interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
@@ 1. Introduction
Home confinement, defined as the restriction of movements within the home (52) due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives (38; 11). Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression (6; 66). For instance, Bartoszek et al. (6) found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue. Understanding these impacts is essential for developing effective targeted interventions to support individuals during such periods.
Traditionally, mental well-being during home confinement has been assessed using survey-based instruments (66; 22), which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias (23; 59). Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative (61; 15). While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation (61; 15), they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited (61). This study shifts focus to more granular data, such as app usage patterns, to provide deeper insights into emotional well-being during confinement.
The diversity of individual circumstances complicates the study of home confinement. Factors such as the reason for confinement, personal health conditions, and social environment can vary greatly, complicating the creation of a standardized model. However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies (57; 70). This shared experience allows for a more uniform analysis of how mobile device usage reflects emotional states, social roles, and work-life dynamics during such periods. Espeically, while this study uses the lockdown as a descriptive case study (87), the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles, during COVID-19 lockdown, as a case study for home confinement. We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health (17; 54). By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect (67) and Social Role from the Role Theory (8), we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics. Our research seeks to address the following questions: RQ1. How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns? RQ2. Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?
@@
To answer these questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia, involving 32 participants. We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) (7) and end-of-day surveys. We also proposed a novel method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records by identifying segments and transition times between apps during the day. The resulting insights offer a deeper understanding of how app usage during confinement reflects emotional well-being and social dynamics, with implications for developing digital interventions to support individuals in long-term home confinement. All scripts for this study are open source. [^1] Link will be shared after the double-blind review. In summary, our contributions are as follows:
@@
  • We collected a rich dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia. This dataset includes 502,485 records of users smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys. It provides a unique, real-world snapshot of emotional well-being and social dynamics during lockdown.
@@
  • We analyzed perceived valence, arousal, and social roles, revealing that the lockdown significantly impacted well-being, reducing emotional valence and increasing work hours due to interruptions or blurred work-life boundaries. These findings shed light on the complex emotional landscape during prolonged confinement.
  • We introduced the App-based Information Gain Temporal Segmentation (AIGTS) method, an unsupervised technique for extracting user activities (e.g., video-watching, socializing, deep work) from app usage data. This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviours relate to emotional well-being.
  • We demonstrated that mobile usage patterns can predict emotional well-being during home confinement. Our extensive experiments revealed key features that are most predictive of emotional states, offering new avenues for developing digital interventions to support individuals during isolation.
@@ 2. Related work
2.1. Impact of the Home Confinement on Mental Well-Being
Home confinement, whether due to edical conditions, disability, or societal events like pandemics, can significantly affect mental well-being. Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness (76; 14; 10). The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger (64; 39). These effects may persist even after restrictions are lifted (31).
The COVID-19 pandemic underscored these effects on a global scale. The World Health Organization (84) reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks. Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread (85; 16; 37). For example, Khubchandani et al. (36) found depression (39%) and anxiety (42%) surged in the U.S. post-lockdown. Similarly, Fountoulakis et al. (20) reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece. Loneliness was also a major issue, especially among those in long-term isolation (37).
Despite extensive research on the mental well-being impacts of various forms of home confinement, much of it is geographically limited or reflects different phases of isolation. This variability makes it challenging to fully understand the broader, long-term psychological effects of home confinement in diverse contexts. To address this gap, RQ1 examines how emotions and social roles are experienced during COVID-19 home confinement, using self-reports and app usage patterns to infer these dynamics.
2.2. Digital Device Usage During Home Confinement
During periods of home confinement, digital devices play a crucial role in maintaining communication, accessing information, and fulfilling work and social roles. With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities (32). For instance, Ratan et al. (62) found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behaviour under prolonged confinement. With workplaces and schools shifting to remote environments, the demand for digital communication tools surged (79; 19; 5). Video conferencing platforms, social media, and messaging apps saw massive increases in usage (70; 57), and overall digital device use skyrocketed. A global survey found 70% of internet users reported increased smartphone usage during the lockdown (70), emphasizing how integral smartphones became in managing both work and personal life. Similarly, Majumdar et al. (45) observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown. Jonnatan et al. (33) also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication (43; 29). Studies suggest that behaviours such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase (46). This highlights the importance of understanding digital device usage patterns in today’s technology-reliant environments, particularly in contexts of home confinement.
@@ 2.2. Digital Device Usage During Home Confinement
2.3. Inferring Mental Well-Being Using Mobile Sensing Technologies
Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood (49; 88; 48), engagement (24; 28; 2), and mental health issues like depression and anxiety (15; 61; 81). Many studies have integrated data from multiple sensors to improve prediction accuracy. For example, Gao et al. (24) used wearable and environmental sensors to predict student engagement, while Wang et al. (81) tracked depression using passive mobile sensing and wearables. Jaques et al. (30) modelled student mood using physiological signals and mobile behaviour.
While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviours like physical activity, communication patterns, and location (25). Studies have shown that mobile usage data can predict psychological states such as anxiety (61; 77), mood (88; 48), and emotion (75). For instance, Zhang et al. (88) developed an app to detect compound emotions from smartphone usage, and Tlachac et al. (77) used call and text logs to predict anxiety and depression during the pandemic.
Research on app sequences has also provided insights into daily behaviours, such as social roles (3) and interruptibility (55; 56). Anderson et al. (3) demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al. (55; 56) used app sequences to predict activity breakpoints for intelligent notification delivery. However, most studies have focused on identifying breakpoints rather than modeling user activity or predicting contextual variables. Additionally, earlier studies often relied on basic aggregation methods, which fail to capture transitions between activities. Thus, RQ1 also explores emotions and social roles by analyzing app usage patterns.
@@ 2.3. Inferring Mental Well-Being Using Mobile Sensing Technologies
While mobile sensing technologies have shown potential in predicting well-being, gaps remain in understanding how these tools can be applied specifically during home confinement. Home confinement introduces unique behaviours—such as increased reliance on digital devices and altered perceptions of time—that differ from daily life. Although some studies (77; 53; 74) have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement. Understanding this unique setting could provide valuable insights for other confinement scenarios, such as post-operative recovery, parental leave, or aging. Therefore, RQ2 investigates how mobile sensing can model emotions during home confinement and identify key features for prediction.
3. Study Overview
We conducted a three-week field study in Melbourne, Australia, from 24 August 2020 to 13 September 2020, during the strictest Stage IV COVID-19 restrictions. This study use Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work. The following section details the background, participants, and data collected.
3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia